OpenStudy (haileyd):

What is the largest decimal number that can be represented by a binary number with 4 place values? (I know each place in a binary number has a value of a power of 2, starting in the ones place with 20.) But yeah i forgot..

OpenStudy (rsmith6559):

1 * 2^0 = 1 1 * 2^1 = 2 1 * 2^2 = 4 1 * 2^3 = 8 8 + 4 + 2 + 1 = 15

OpenStudy (haileyd):

Thanks